OpenStudy (anonymous):

Help??? In your own words, explain how a trinomial of the form 2x2 + 13x + 15 can be turned into a four term polynomial suitable for factoring by grouping. Use complete sentences.

OpenStudy (callisto):

2x2 + 13x + 15 = 2x^2 + 10x +3x +15 = 2x(x+5) + 3(x+5) = (2x+3)(x+5)